DX LUCI/AO Commissioning

01:00 ACE Preset to 10 mag Reference Star
  • Not seen on CCD47
  • Steve had checked pointing 2 degrees away
  • Steve checked again. It was close
  • Next ACE Preset had not AO Reference Star
  • Next field

01:22 ACE Preset to 12.9 mag
  • See 13.5
  • Bin 3, 200 Hz
  • Loop Closed

01:30 HUMAN ERROR: Clicked Cancel Preset on FLAO GUI rather than Done
  • Closed Loop manually but could not get AOS to see that the AO Loop is closed
  • Tried Pause, Resume and StartAO from AOS GUI. Get incorrect arbitrator state "Ready"
  • Also received setGain error
AdSecArbitrator_292|INF|       247|2017-11-02 01:32:08.972181|   COMMANDHANDLER > FSM (status: AOSet) has received command 2018 (SetGain)
20170211 01:32:08.9734 Alert: Warning: setGain not executed (retry)
AdSecArbitrator_292|WAR|       248|2017-11-02 01:32:08.972930|              IDL > Command returns retryable error : IDL File error
AdSecArbitrator_292|INF|       249|2017-11-02 01:32:09.064982|   COMMANDHANDLER > Command not executed (status RETRY) (errstr IDL File error)
  • the gain0.fits file was not found: /local/aomeas/adsec_calib/M2C/KL_v18/gain/gain0.fits
  • % cp /local/aomeas/adsec_calib/M2C/KL_v20/gain/gain0.fits /local/aomeas/adsec_calib/M2C/KL_v18/gain/gain0.fits

01:40 Send new preset from IDL, LEFT only
  • iif_preset, CURRENT_LEFT
    • Successful ACE preset with the same parameters as the previous LUCI preset
  • iif_runao
    • Successful

01:48 Loop close on LEFT (both in fact)
  • ready for observations
  • LUCI SW problem with AFC moving Fold Mirrors to align mask

02:48 LUCI working again

02:53 Actuator 231 has a high current, -0.25 N
  • This value has been high since the beginning of the night
  • Only had 241 modes in flat that usually has 650 modes

04:10 - 04:11 DX camera lens run away. Loop open during runAO. Manually set the camera lens position and re-closed the loop. Loop was then stable.
  • This happens occasionally in bin 4 on DX. SX does not seem to have this problem.

05:00ish ACE Preset
  • 14.4 mag catalog
  • 15.2 mag measured
  • bin 4, 200 Hz

05:56 - 05:57 SX Skip Frames, open loop, pause LUCI, close loop, continue LUCI
  • Seeing is changing from 0.9 to 1.2

06:27 - 06:28 DX Skip Frames, open loop, pause LUCI, close loop, continue LUCI
  • Seeing is changing from 0.9 to 1.2
  • Actuators on along the bottom edge with high forces
    • Turned down gains (2.4, 0.9, 0.9 => 1.9, 0.9, 0.9)

06:33 - 06:35 SX Skip Frames, open loop, pause LUCI, close loop, continue LUCI
  • IIF indicated the SX Shell had RIPPed. It had not.
  • Guiding was turned back on. Pulled star off the tip of the pyramid
  • Had to change Rx by -6" to get light back on pyramid
  • Then close loop

07:10 LUCI/AO Commissioning complete. Hand over to LBTB

-- %USERSIG{DougMiller - 2017-11-02}%
Topic revision: r2 - 02 Nov 2017, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback